For given values of the sample mean and the sample standard deviation when n = 25, you conduct a hypothesis test and obtain a p-value of 0.0667, which leads to non-rejection of the null hypothesis. What will happen to the p-value if the sample size increases (and all else stays the same)?

A. Increase
B. Stay the same
C. Decrease
E. May either increase or decrease
1
Expert's answer
2020-06-10T19:56:02-0400

Use a t-statistic, with degrees of freedom "df=n\u22121"


 for inference for a population mean:


"CI: x\u00af\u00b1t^{\u2217}_{df}SE" "HT: Tdf\frac{x¯−μ}{SE}"


"where SE=s/\sqrt n"


Use a t-statistic, with degrees of freedom "df=n_{diff}\u22121"  for inference for the difference in two paired (dependent) means:

"CI: x_{diff}\u00b1t^\u2217_{df}\/SE"


"SE=s/\sqrt n" Note that "\u03bc_{diff}" is often 0, since often "H_{0}:\u03bc_{diff}=0 ."


As the sample size increases the standard error will decrease, which increases the test statistic, and hence decreases the p-value